# Deep Generative Model for Human Mobility Behavior

### v1

@@ Methods
We focus on conditional mobility generation, aiming to simulate a sequence of future activity events $\mathbf{y}=\left[y_{1},y_{2},...,y_{k_{1}}\right]$ based on a preceding sequence of observed events $\mathbf{x}=\left[x_{1},x_{2},...,x_{k_{2}}\right]$ by the same individual. Each activity event $x_{i}$ or $y_{i}$ is represented as a tuple $\langle l,t,d,m\rangle$, where $l$ denotes the location identifier, $t$ the start time, $d$ the duration, and $m$ the travel mode. To incorporate spatial context, each event in the observed sequence $\mathbf{x}$ is further associated with geographic coordinates $g$ and land-use context $c$ describing the surrounding environment.
Data sources and preprocessing
Mobility Data
We selected 2,113 individuals who had been tracked for at least 50 days and whose locations were known at least 60% of the time. This resulted in a dataset with an average tracking duration of 138 days and an average temporal coverage of 88% across individuals.
Extracting activity attributes
External contextual data
We collected the point of interest (POI) dataset from OpenStreetMap (OSM) (http://www.openstreetmap.org), an open-source project that provides free and easily accessible digital map resources. Historical POI data for Switzerland were acquired from the end of 2022 to align with the time frame of the tracking dataset. The POI dataset was supplemented with OSM building data from the same year. For each building entry, we retained its attribute values but created a new POI by using the center point of its polygon geometry as its spatial representation. After removing redundant and insignificant POIs, the final dataset consisted of 3,104,590 POIs, categorized into 19 first-level and 188 second-level groups. The second-level categories, representing functional type descriptions, were used to obtain urban functions for locations.
Data splits
We divided the activity event sequences into non-overlapping training, validation, and test sets in a 7:2:1 ratio based on time. Sequences occurring in the first 70% of tracking days for each individual were designated as training data, while the last 10% of days were used as test data. We considered mobility over the past three weeks as the traveled sequence and activity events in the subsequent week as the target sequence. The lengths of these sequences, as denoted by $k_{1}$ and $k_{2}$, vary depending on the specific data. This process yielded 590,257, 104,010, and 36,285 pairs of traveled-target sequences for training, validation, and testing, respectively.
Diffusion-based modeling of discrete activity sequences
@@
To control the generated data explicitly, we can introduce a guidance signal $x$ to learn the conditional distribution $q(\mathbf{z}\mid x)$. A natural way is to supplement the conditioning information at each transition step during the original reverse process (Equation 3):
$p_{\theta}(\mathbf{z}_{0},...,\mathbf{z}_{T}\mid x)\coloneq p(\mathbf{z}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},x)$ (8)
The unconditional model $\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t)$ is replaced with the conditional model $\hat{\mathbf{z}}_{\theta}(\mathbf{z}_{t},t,x)$, while the diffusion process remains unchanged. In MobilityGen, the guidance signal is derived from the traveled mobility events $\mathbf{x}$, and the model learns to generate a new target sequence $\mathbf{y}$ based on this signal (Figure 1).
Neural network architecture
The neural network comprises several components: an embedding module that maps activity attributes into an embedding space, an encoder that extracts information from traveled traces, a decoder that reconstructs target sequences from their corrupted versions, and task-specific linear heads that ensure consistency mapping.
Inputs and embedding modules
@@
For each record $x_{k}$ in the traveled sequence, we also incorporate the geometry $g_{k}$ and surrounding land-use functions $c_{k}$. The geometry is modeled using the Space2vec method [Mai et al., 2020], which transforms raw coordinates in a projected coordinate system into vector representations via sine and cosine base functions. Land-use functions are captured using the location-POI embedding method proposed by Hong et al. [2023], which considers POIs from various categories within a location and applies Latent Dirichlet Allocation (LDA) to derive their functional descriptions. The relationships between these movement-related contexts and the activity embeddings are learned through feed-forward networks $f$ and linearly combined via a residual connection:
@@
Encoder and decoder design
We utilize the transformer encoder-decoder architecture proposed by Vaswani et al. [2017]. The encoder processes the traveled sequence representation through $L$ encoder blocks, each with an identical architecture consisting of a multi-head self-attention layer and a feed-forward layer with two linear operations separated by a non-linear activation function. Residual connections [He et al., 2016], layer normalizations [Ba et al., 2016], and dropout [Hinton et al., 2012] are applied to each layer. The decoder, which processes the corrupted target sequence representation, contains $L$ decoder blocks structured similarly to the encoder. In addition to the two layers in each encoder block, the decoder incorporates an intermediate layer that performs multi-head attention over the encoder stack’s output.
@@
where $\alpha$ values are weights that balance the contributions of the individual losses.
Training procedure and computational cost
@@
Evaluation metrics for mobility modeling
We use metrics that summarize high-level mobility patterns at collective and individual levels to quantify the discrepancy between ground truth and generated activity sequences. Here, we consider the following indicators:
@@
    Location visitation frequency [González et al., 2008, Song et al., 2010]. Individuals’ preferences in location choices are highly unbalanced and can be quantified by the relationship between the visitation frequency of locations $f_{k}$ and their rank $k$, which is reported to follow Zipf’s law, i.e., $f_{k}\sim k^{-\xi}$, with $\xi\sim 1.2$.
@@
    Temporal mobility entropy [Song et al., 2010]. Closely related to the theoretical predictability of location sequences, temporal mobility entropy $S_{temp}$ captures the full spatio-temporal order of location visits by measuring the probability $P(T_{u}^{{}^{\prime}})$ of observing ordered subsequences $T_{u}^{{}^{\prime}}$ within the location sequence $T_{u}$:
    $S_{temp}=-\sum_{T_{u}^{{}^{\prime}}\in T_{u}}P(T_{u}^{{}^{\prime}})\log_{2}(P(T_{u}^{{}^{\prime}}))$
    We use the method described in Song et al. [2010] to estimate the temporal mobility entropy $S_{temp}$ for individuals.
@@
    Evolution of trip package choices [Hong et al., 2023]. The trip package concept was introduced to encapsulate the interactions within multifaceted activity-travel patterns. For each individual, a trip package is constructed when observing the usage of a travel mode to reach an activity location, i.e., a unique mode-location choice. Individuals continuously explore new activity-behavior options, and the number of unique trip packages observed over time quantifies the speed of this exploratory behavior.
Identifying urban functions from POI features
POIs reflect the socioeconomic activities in local regions and have been increasingly used to identify urban functions at a fine scale [Yao et al., 2017]. A well-established approach to functional identification involves obtaining a POI representation for each study unit, identifying cluster structures from these representations, and labeling the resulting classes based on expert knowledge [Niu and Silva, 2021]. Here, we perform $k$-means clustering on the LDA functional descriptor of each location. Using the silhouette score to assess clustering quality, we determine the optimal number of classes to be five (Extended Data Figure 3a). Visualizing the five-class clustering result using 2-dimensional t-SNE reveals that the data points within each class are compact, while those between classes are well-separated (Extended Data Figure 3b). To interpret the urban functions of each location group, we rely on the spatial distribution of classes (Extended Data Figure 3c) and the POI enrichment factor for each class (Extended Data Table 5). The latter quantifies the relative occurrence of different categories of POIs. For each location class $i$ and POI category $q$, the enrichment factor $EF_{i}^{q}$ is defined as:
$EF_{i}^{q}=\frac{N_{i}^{q}/N_{i}}{N^{q}/N}$
Deriving experienced income segregation
To measure income segregation, we divide individuals into four roughly equal-sized quartiles according to their reported monthly household income: low (<4,000<4,000 CHF), lower-middle (4,0008,0004,000-8,000 CHF), upper-middle (8,00012,0008,000-12,000 CHF), and high (>12,000>12,000 CHF). We then apply the experienced income segregation framework introduced by Moro et al. [2021], described in detail below.
$S_{u}=\frac{2}{3}\sum_{q}\left|\tau_{uq}-\frac{1}{4}\right|$

### v2

@@ 1.2 Problem definition
Formally, we focus on conditional mobility generation, aiming to simulate a sequence of future activity events $\mathbf{y}=\left[y_{1},y_{2},...,y_{k_{1}}\right]$ based on a preceding sequence of observed events $\mathbf{x}=\left[x_{1},x_{2},...,x_{k_{2}}\right]$ by the same individual. Each activity event $x_{i}$ or $y_{i}$ is represented as a tuple $\langle l,t,d,m\rangle$, where $l$ denotes the location identifier, $t$ the start time, $d$ the duration, and $m$ the travel mode. To incorporate spatial context, each event in the observed sequence $\mathbf{x}$ is further associated with geographic coordinates $g$ and land-use context $c$ describing the surrounding environment.
1.3 Diffusion-based modeling of discrete activity sequences
@@
1.4 Neural network architecture
MobilityGen’s neural network comprises several components: an embedding module that maps activity attributes into an embedding space, an encoder that extracts information from traveled traces, a decoder that reconstructs target sequences from their corrupted versions, and task-specific linear heads that ensure consistency mapping.
@@ 1.4 Neural network architecture
For each record $x_{k}$ in the traveled sequence, we also incorporate the geometry $g_{k}$ and surrounding land-use functions $c_{k}$. The geometry is modeled using the Space2vec method [Mai et al., 2020], which transforms raw coordinates in a projected coordinate system into vector representations via sine and cosine base functions. Land-use functions are captured using the location-POI embedding method proposed by Hong et al. [2023b], which considers POIs from various categories within a location and applies Latent Dirichlet Allocation (LDA) to derive their functional descriptions. The relationships between these movement-related contexts and the activity embeddings are learned through feed-forward networks $f$ and linearly combined via a residual connection:
@@
We utilize the transformer encoder-decoder architecture proposed by Vaswani et al. [2017] as the network backbone. The encoder processes the traveled sequence representation through $L$ encoder blocks, each with an identical architecture consisting of a multi-head self-attention layer and a feed-forward layer with two linear operations separated by a non-linear activation function. Residual connections [He et al., 2016], layer normalizations [Ba et al., 2016], and dropout [Hinton et al., 2012] are applied to each layer. The decoder, which processes the corrupted target sequence representation, contains $L$ decoder blocks structured similarly to the encoder. In addition to the two layers in each encoder block, the decoder incorporates an intermediate layer that performs multi-head attention over the encoder stack’s output.
@@
where $\alpha$ values are weights that balance the contributions of the losses.
2 Experiment
2.1 Data sources and preprocessing
We selected 2,113 individuals who had been tracked for at least 50 days and whose locations were known at least 60% of the time. This resulted in a dataset with an average tracking duration of 138 days and an average temporal coverage of 88% across individuals. The dataset provides substantial longitudinal depth and multi-attribute annotation across diverse Swiss contexts, supporting learning of the coupled spatio-temporal and modal structure of daily mobility.
For acquiring the built environment information, we collected the POI dataset from OpenStreetMap (OSM) (http://www.openstreetmap.org), an open-source project that provides free and easily accessible digital map resources. Historical POI data for Switzerland were acquired from the end of 2022 to align with the time frame of the tracking dataset. The POI dataset was supplemented with OSM building data from the same year. For each building entry, we retained its attribute values but created a new POI by using the center point of its polygon geometry as its spatial representation. After removing redundant and insignificant POIs, the final dataset consisted of 3,104,590 POIs, categorized into 19 first-level and 188 second-level groups. The second-level categories, representing functional type descriptions, were used to obtain urban functions for locations.
We divided the activity event sequences into non-overlapping training, validation, and test sets in a 7:2:1 ratio based on time. Sequences occurring in the first 70% of tracking days for each individual were designated as training data, while the last 10% of days were used as test data. We considered mobility over the past three weeks as the traveled sequence and activity events in the subsequent week as the target sequence. The lengths of these sequences, as denoted by $k_{1}$ and $k_{2}$, vary depending on the specific data. This process yielded 590,257, 104,010, and 36,285 pairs of traveled-target sequences for training, validation, and testing, respectively. This provides a large set of conditional sequence pairs for multi-attribute schedule generation and supports stable learning in our experiments.
2.2 Training procedure and computational cost
@@
2.3 Evaluation metrics
We use metrics that summarize high-level mobility patterns at collective and individual levels to quantify the discrepancy between ground truth and generated activity sequences. We consider the following indicators:
@@ 2.3 Evaluation metrics
    Location visitation frequency [González et al., 2008, Song et al., 2010b]. Individuals’ preferences in location choices are highly unbalanced and can be quantified by the relationship between the visitation frequency of locations $f_{k}$ and their rank $k$, which is reported to follow Zipf’s law, i.e., $f_{k}\sim k^{-\xi}$, with $\xi\sim 1.2$.
@@
    Temporal mobility entropy [Song et al., 2010a]. Closely related to the theoretical predictability of location sequences, temporal mobility entropy $S_{temp}$ captures the full spatio-temporal order of location visits by measuring the probability $P(T_{u}^{{}^{\prime}})$ of observing ordered subsequences $T_{u}^{{}^{\prime}}$ within the location sequence $T_{u}$:
    $S_{temp}=-\sum_{T_{u}^{{}^{\prime}}\in T_{u}}P(T_{u}^{{}^{\prime}})\log_{2}(P(T_{u}^{{}^{\prime}}))$ (19)
    We use the method described in Song et al. [2010a] to estimate the temporal mobility entropy $S_{temp}$ for individuals.
@@
    Evolution of trip package choices [Hong et al., 2023a]. The trip package concept was introduced to encapsulate the interactions within multifaceted activity-travel patterns. For each individual, a trip package is constructed when observing the usage of a travel mode to reach an activity location, i.e., a unique mode-location choice. Individuals continuously explore new activity-behavior options, and the number of unique trip packages observed over time quantifies the speed of this exploratory behavior.
3 Results
3.1 Microscopic modeling reproduces location choice behavior
To test whether our event-level formulation preserves the core spatial regularities of daily mobility, we begin by evaluating MobilityGen’s ability to capture location patterns. How individuals select spatial locations for different activities shapes their interactions with physical space and is a fundamental driver of many urbanization-induced challenges [Moro et al., 2021, Nilforoshan et al., 2023, Liu et al., 2023]. Here, we focus on the simulated visits to locations and compare them with the real visits from observed sequences.
We first assess the ability of MobilityGen to capture fundamental properties of location choices and compare its performance against two state-of-the-art microscopic mobility models, namely, exploration and preferential return (EPR) [Song et al., 2010b], and Container [Alessandretti et al., 2020] (Figure 2). We begin with the average rank distribution of location visits (Figure 2a), a widely used measure of preference for activity locations [González et al., 2008, Barbosa et al., 2018]. Our results indicate that MobilityGen provides a significantly better fit to real data than both the EPR and Container models, as confirmed by the likelihood ratio test [Clauset et al., 2009] (with $P<10^{-3}$) and the Wasserstein distances [Panaretos and Zemel, 2019] between the rank distributions (see Appendix Supplementary Note 1.1 and Table S1). The radius of gyration, which measures the characteristic distance an individual covers [González et al., 2008], serves to quantify the movement intensity in space. We observe that the median evolution of the radius of gyration across individuals, $r_{u}(k)$, over displacement steps $k$, is well described by logarithmic growth for both real and generated location sequences (Figure 2b), consistent with previous findings [Alessandretti et al., 2020]. However, when fitting $r_{u}(k)\propto\alpha\log(k)$, only the value $\alpha$ derived from MobilityGen agrees well with the real data (see Appendix Supplementary Note 1.1). Finally, we evaluate the temporal mobility entropy that quantifies the predictability of location sequences, considering both the frequency and the order of visits [Song et al., 2010a]. Using the likelihood ratio test and the Wasserstein metric, we further confirm that, compared against the EPR and Container models, MobilityGen most accurately approximates the true entropy distribution over the population (with $P<10^{-3}$, Figure 2c and see Appendix Supplementary Note 1.1 and Table S1). Additionally, we evaluated a representative set of state-of-the-art data-driven mobility models, including deep learning-based sequence prediction models and end-to-end generative models, and further location metrics (see Appendix Supplementary Note 1.1). Our findings indicate that MobilityGen consistently provides the best fit to real-world data. Evaluations with metrics from language modeling confirm that MobilityGen can simulate location visits that closely resemble the real sequence while avoiding regurgitation of input sequences and maintaining sufficient diversity (see Appendix Supplementary Note 1.2).
To further evaluate the generated mobility patterns, we examine the spatial distribution of location visits. As an illustrative example, we map the observed and synthetic trajectories of a sample individual (Figure 3a and b). The comparison shows that MobilityGen effectively captures both visit and transition patterns at frequently visited locations (thick lines), while also generating occasional visits to less familiar locations (thin lines). Accurately modeling such exploratory behavior remains a central challenge in human mobility research [Cuttone et al., 2018]. To assess how well MobilityGen addresses this challenge at the population level, we compare the aggregated spatial distribution of location visits across Switzerland in the real and simulated data (Figure 3c and d). Visit frequencies are color-coded from low (yellow) to high (red). MobilityGen closely matches the concentration of frequently visited locations and also generates visits to less frequented areas, reflecting a balance between individual-level exploration and population-level spatial preference. Using the entropy $E=-\sum_{k}p(f_{k})\log p(f_{k})$ to quantify the dispersion of location visits $f_{k}$, we obtain 0.33 for the real data and 0.53 for the simulated data, indicating greater spatial dispersion in the latter. To analyze spatial dynamics at a finer resolution, we examine three representative cities: Bern, Zurich, and Lucerne (Figure 3e-j), each characterized by distinct urban functions and mobility patterns. MobilityGen successfully reproduces key urban visitation structures in all three cases. However, spatial dispersion remains higher in the simulated data, with entropy values of 2.49 vs. 1.89 for Bern, 3.99 vs. 3.39 for Zurich, and 2.71 vs. 2.11 for Lucerne. Despite this discrepancy, MobilityGen compares favorably to other microscopic mobility models in replicating the location visit distributions (see Appendix Supplementary Note 1.1, Table S2). Complementary visual comparisons are provided at both national and urban levels in Appendix Supplementary Note 1.1, Figure S2 and S3.
3.2 Structured behavioral patterns across time, mode, and location
To examine the proposed coupled representation beyond location patterns, we evaluate whether MobilityGen jointly captures activity timing, travel mode, and their interdependencies within unified activity events. We assess these structured patterns for each sequence in the test set and compare cross-attribute regularities and interactions (Figure 4).
Incorporating temporal information allows MobilityGen to reason about the scheduling of each activity. This temporal dimension is evident in the activity durations (Figure 4a) and the number of daily visited locations (Figure 4b), both of which closely match those observed in the real data. To further examine patterns of activity participation, we construct mobility motifs [Schneider et al., 2013], which serve as high-level descriptors of regularities in activity location and scheduling. MobilityGen accurately recovers the relative frequencies of both common and rare motifs, including simple two-location transitions and more complex patterns involving up to five locations (Figure 4c). We further benchmark MobilityGen against two state-of-the-art spatio-temporal mobility simulators, the Diary-based Trajectory Simulator (DITRAS) [Pappalardo and Simini, 2018] and TimeGeo [Jiang et al., 2016], using additional mobility metrics (see Appendix Supplementary Note 2.1). Across all measures, MobilityGen more faithfully reproduces the interplay between activity timing and location choice, yielding the closest alignment with empirical behavior.
Next, we incorporate user-verified travel mode labels into MobilityGen to jointly model mode choice in relation to activity events. Figure 4d compares the travel mode distributions from the observed and simulated activity sequences, with discrepancies quantified by the Wasserstein distance (see Appendix Supplementary Note 2.1, Table S4). Across all scenarios, MobilityGen reconstructs the mode preferences exhibited in real-world behavior. Car and walking, the two most common modes, show excellent agreement between empirical and simulated shares. Notably, MobilityGen captures subtler distinctions among public modes, including bus, train, and tram, accurately reflecting their relative frequencies. We further assess mode-specific behavior by comparing the displacement distributions for car and walking trips (Figure 4e and f). For car travel, the simulated displacements closely align with the observed distribution (Kolmogorov–Smirnov statistic $KS=0.02$, with $P<10^{-3}$). For walking, MobilityGen produces slightly longer displacements ($KS=0.06$, $P<10^{-3}$), potentially due to short trips being disproportionately affected by spatial discretization. This effect could be mitigated in future work by adopting a finer location grid, although such refinement may increase computational complexity.
Finally, we assess the evolution of trip packages, defined as combinations of travel modes and activity locations [Hong et al., 2023a], to evaluate MobilityGen’s ability to model complex attribute interactions (Figure 4g). The average number of unique trip packages over time, denoted as $Tp(t)$, follows a sub-linear growth over time $t$, consistent with previous findings [Hong et al., 2023a]. The scaling exponent $\alpha$ in the relation $Tp(t)\propto t^{\alpha}$ is nearly identical between simulated ($\alpha=0.85\pm 0.01$) and empirical sequences ($\alpha=0.85\pm 0.01$), indicating that MobilityGen accurately captures the evolving diversity of travel mode and activity location choices over time. Further ablation studies confirm that the full multi-attribute model achieves a more realistic representation of attribute interactions than simplified variants with fewer modeled attributes (see Appendix Supplementary Note 3.2).
To understand how MobilityGen captures attribute interactions, we analyze the structure of its learned embedding space. We apply densMAP [Narayan et al., 2021] to project the 128-dimensional embeddings into two dimensions for visualization (see Appendix Supplementary Note 2.2). The projections reveal that MobilityGen learns semantically meaningful representations: similar travel modes appear near each other, temporal attributes vary smoothly along continuous trajectories, and frequently visited locations are well-separated from infrequent ones. When multiple attributes are jointly embedded, the space further reveals distinct event clusters tied to specific locations, alongside a more diffuse region that likely reflects less location-dependent activities. Together, these patterns indicate that the embedding space offers a compact representation of multi-attribute mobility events and a natural basis for quantitative analysis of structure across mode, time, and location. By providing this shared representation, MobilityGen enables complex mobility dynamics to be inferred directly from behavioral sequences and facilitates diverse mobility analyses.
3.3 Context enhances generalization to novel locations
A key element of our formulation is that contextual information can influence individual travel decisions – an aspect well-supported by the literature but not adequately covered by existing microscopic mobility models [Pappalardo et al., 2023]. In MobilityGen, context features are combined with activity attributes through learnable representations, yielding a context-aware mobility model (see Figure 1b). We focus on spatial contexts that characterize the physical environment, namely coordinate geometry, which reveals movement directionality [Zhao et al., 2024]; and POIs, which display urban functions and imply movement purposes [Hong et al., 2023b]. Further contexts could be integrated through appropriate representations; we defer this to future research.
We analyze location patterns to assess the influence of contextual factors on the simulation performance. Generated locations are compared to observed data and categorized into three sets: Set A includes locations present in both sequences, while Sets B and C contain locations visited exclusively in the observed and generated sequences (Figure 5a). Since MobilityGen replicates key location preferences (see Figure 3), Set A is expected to capture the locations where individuals conduct most of their activities. Indeed, an analysis of the visit frequencies reveals that 96.4% of the true visits and 93.4% of the generated visits occur at locations in Set A (Figure 5b). In contrast, locations in Sets B and C receive relatively few visits, reflecting individuals’ tendencies toward novelty-seeking and the exploration of unfamiliar locations [Arentze and Timmermans, 2005, Song et al., 2010b, Alessandretti et al., 2018].
To understand how MobilityGen models novelty-seeking behavior, we identify the functional types of locations using POI data. Applying $k$-means clustering to POI representations of locations yields five distinct location classes, whose functional types are inferred from their spatial distributions and the enrichment of specific POI categories [Yao et al., 2017, Niu and Silva, 2021] (see Appendix Supplementary Note 3.1). This functional classification enables a detailed analysis of location preferences, revealing that visits in Set A predominantly occur in the City center, whereas visits in Sets B and C are largely concentrated in Residential & Countryside. Moreover, preferences for Leisure & Recreation and Nature & Tourism sites are markedly higher in Sets B and C than in Set A. Using pairwise Wasserstein distances to quantify the similarity of visitation patterns across sets, we find that Sets B and C are closely aligned (Wasserstein distance: 0.11), while their distributions differ substantially from that of Set A (Set A vs. Set B: 0.57; Set A vs. Set C: 0.56). In other words, MobilityGen captures key patterns of exploratory behavior, despite the inherent difficulty of characterizing novel locations and predicting which specific places individuals are likely to explore.
A comparison with a model that excludes contextual information confirms that incorporating context improves the ability to characterize locations (see Appendix Supplementary Note 3.2). This improvement is particularly relevant in settings where historical mobility information is limited, such as when individuals seek novel locations for their activities. To benchmark how effectively MobilityGen leverages contextual cues, we conducted comparative experiments against established mobility flow generation models: Gravity [Zipf, 1946], Deep Gravity [Simini et al., 2021], and random forest [Cabanas-Tirapu et al., 2025]. These baselines estimate flows between location pairs using features such as distance, population, and selected facility densities (e.g., shops and transport hubs). Flow-based models differ fundamentally from MobilityGen in both problem formulation and modeling capacity. For instance, they cannot capture temporal dynamics or disaggregated mobility. To ensure a fair comparison, we focused exclusively on locations that were not observed during training. The results show that MobilityGen outperforms the flow generation baselines by a substantial margin across multiple flow metrics (see Appendix Supplementary Note 3.3 and Table S9). We attribute this improved performance in novel locations to the integration of spatial features, including coordinate geometry and detailed POI information, as well as the joint modeling of contextual and behavioral factors. While further research is needed to fully understand novelty-seeking mobility, our findings support the hypothesis that comprehensive context modeling is necessary to capture the variability in human behavior.
3.4 Reproducing mode-specific spatial behavior and co-presence dynamics
Beyond standard validation, we ask whether the proposed event-level representation enables analyses that have so far been difficult or infeasible with existing simulators. Our simulations generate realistic daily schedules that jointly capture where people go, how they travel, and when activities occur. Earlier tools often rely on predefined rules or cover only part of this behavioral structure, which has limited their use for studying differences across travel modes or patterns of shared presence in everyday settings. We focus on two representative scenarios with broad relevance to transport planning and urban research: (i) mode-specific spatial visitation, showing how individuals use different travel modes to move through urban space, and (ii) spatio-temporal co-presence, reflecting how individuals co-occur in space and time within shared environments.
While spatial visitation captures where individuals go, co-presence patterns reveal whom they encounter, offering complementary insight into the intensity and distribution of interpersonal contact in daily urban life. A key application of co-presence lies in socio-spatial segregation research, where attention has shifted from static residential indicators to individual-level experienced segregation enabled by fine-grained mobility data [Liao et al., 2025]. We assess MobilityGen’s ability to reproduce these patterns by focusing on income-based segregation. We group individuals into four income quantiles and measure, at each location, the time spent by people from different income groups [Moro et al., 2021] (see Appendix Supplementary Note 4). These measures are then combined into an individual-level segregation score that reflects each person’s exposure to others across income levels. Simulated and observed segregation scores show strong agreement (Pearson $r=0.70$, $P<10^{-3}$; Figure 7), capturing both overall trends and the variability in co-presence-based income segregation. A breakdown by income quantiles confirms that MobilityGen reproduces systematic differences in exposure and substantially outperforms baseline simulators such as TimeGeo and DITRAS, whose separation of space and time limits their ability to capture realistic contact patterns (see Appendix Supplementary Note 4, Figure S8). This case study illustrates MobilityGen’s broader value as a generative model of social exposure. By capturing individual-level interactions in daily schedules, it provides a simulation environment to explore how population-level social mixing evolves under different scenarios. This extends beyond segregation analysis to domains such as information diffusion and infectious disease spread, where spatial co-presence fundamentally shapes outcomes.
4 Discussion and conclusion
Despite growing interest in understanding human movement patterns and their implications, existing simulation models fail to capture the full spatio-temporal variability of mobility behavior. Here, we address this gap by advancing a unified event-level formulation of daily mobility that generates sequences of multi-attribute mobility events over days and weeks. We instantiate this formulation in MobilityGen, which provides strong alignment with empirical patterns across key mobility indicators and captures coupled structure across time, space, and mode that earlier approaches struggle to represent. By linking behavioral attributes with contextual information, MobilityGen enables examination of how urban space is accessed differently across travel modes and how co-presence dynamics shape patterns of social exposure, such as income-based segregation. These examples show how generative modeling moves beyond reproducing standard metrics to open new avenues for research and policy enabled by fine-grained mobility data.
Our data-driven framework offers a general event-level formulation for modeling individual mobility behavior. By integrating behavioral attributes into a shared latent space, MobilityGen captures their interactions across extended temporal and spatial scales. This enables examination of complex, longitudinal, and nuanced behavioral patterns that are otherwise difficult to represent using traditional approaches. Looking ahead, MobilityGen could be combined with utility-based activity models in a hybrid approach, improving interpretability and enabling simulations that account for household activity participation and targeted mobility interventions. Its flexibility also supports the integration of richer contextual information, from socio-demographic characteristics to social networks and environmental factors such as weather. Together, these extensions position MobilityGen as a bridge between mobility research and disciplines such as social, health, and environmental sciences, representing a step toward an integrative framework for interdisciplinary mobility modeling.
Data availability
Raw data for the MOBIS dataset are not publicly available due to privacy considerations, but aggregated trip-level data are available to reviewers upon request. OpenStreetMap POIs used for context were downloaded from Geofabrik (region: Europe/Switzerland; snapshot: late 2022) and can be downloaded from https://download.geofabrik.de/.
Code availability
All code to preprocess movement data and train MobilityGen is available at https://github.com/mie-lab/mobility_generation (Apache-2.0 license). Analyses and visualizations were performed in Python and QGIS 3.22.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used GPT-5 in order to improve the language, grammar and readability. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
Appendix